OpenStudy (anonymous):

Javier invests money in two accounts. He invests twice as much in an account paying 11% as he does in an account paying 3%. If he earns $104.50 in interest in one year from both accounts combined, how much did he invest altogether?

OpenStudy (anonymous):



OpenStudy (anonymous):

we have a total interest = a + b

OpenStudy (anonymous):

104.5 = (0.11(2x)) + (0.03(x))

OpenStudy (anonymous):

104.5=.22x+.03x

OpenStudy (anonymous):

.25x=104.5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=418

OpenStudy (anonymous):

2(418)+418 = total investment

OpenStudy (anonymous):

1254

OpenStudy (anonymous):

i believe we got it

OpenStudy (anonymous):

but when i see the answers they are 800$, 550$, 1200$, 1350$

OpenStudy (anonymous):

or none of the above

OpenStudy (anonymous):

hm lets check our math

OpenStudy (anonymous):

i think our math is right

OpenStudy (anonymous):

i believe so as well so it might just be none of the above

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

well thank you for helping me !

OpenStudy (anonymous):

np good luck

OpenStudy (anonymous):

Yes it was correct and thanks!

OpenStudy (anonymous):

^^)